# HyperspectralViTs: General Hyperspectral Models for On-board Remote Sensing

### v2

@@ V-D Inference times on low compute hardware
Table V shows the model inference times needed to process a full EMIT scene capture of 1280x1242 px with our models for methane leak detection. These measurements show the IO time needed to load the hyperspectral datacube and the model inference time - for baselines this also includes the time needed to calculate the MF product.
@@ VI Conclusion
As a potential limitation of the proposed end-to-end HyperspectralViT models, they operate directly on the raw hyperspectral data and as such, they can’t be used for zero-shot generalisation as for example the model of [34]. We however note, that a similar simulated dataset can be created from representative data for any other satellite and later used to train end-to-end models. Secondly, our model doesn’t support additional tasks such as plume quantification or source point estimation. We however note, that the prediction of our model can be used as a binary mask to clean the noisy methane enhancement products, after which classical pipelines for these tasks can be used [62].
For future research, we’d like to explore further steps to make the models more efficient. Our proposed changes to the Transformer based models model can be combined with pruning reported in [63, 38]. Specialised accelerators as those suggested by [64] could also be leveraged. Secondly, we’d also like to explore extension of our models with methods for uncertainty estimation. Methods such as model ensembles [65], or Monte Carlo Dropout [66] could be easily added to the proposed semantic segmentation models, as was previously demonstrated in [67]. Thirdly, we note that the existing archives of EMIT data (currently with more than 107k captures totaling more than 215 TB of data) offer unique opportunities to train large general foundation models on hyperspectral data, where one of the example tested tasks could also be methane detection. Models proposed in this paper could likely be leveraged for general learning tasks (in the direction outlined by our mineral identification task), but also for other downstream applications highlighted by the future planned missions such as NASA’s SBG [58] or ESA’s CHIME [68].
Finally, we are releasing the three newly created datasets on https://huggingface.co/previtus [exact links TBR] and the code and the pre-trained models on Github at https://github.com/previtus [exact link TBR].

### v3

@@ V-D Inference times on low compute hardware
Table V shows the model inference times needed to process a full EMIT scene capture of 1280x1242 px with our models for methane leak detection. We tile this full granule into 100 non-overlapping tiles (padding the data if needed) - we note that in this way we could circumvent limitations of fixed input sizes encountered for example with the Myriad X VPU chip. These measurements show the IO time needed to load the hyperspectral datacube and the model inference time - for baselines this also includes the time needed to calculate the MF product.
@@ VI Conclusion
As a potential limitation of the proposed end-to-end HyperspectralViT models, they operate directly on the bands of hyperspectral data and as such, they can’t be used for zero-shot generalisation as for example the model of [17]. We however note, that a similar simulated dataset can be created from representative data for any other satellite and later used to train end-to-end models. Secondly, our model doesn’t support additional tasks such as plume quantification or source point estimation. We however note, that the prediction of our model can be used as a binary mask to clean the noisy methane enhancement products, after which classical pipelines for these tasks can be used [64].
As a further limitation, we also note, that we used the L1B level of data processing in our experiments instead of the truly “raw” versions of the data (sometimes called L0). This is due to the fact that this data is not available for these sensors, in fact this is a common problem in the wider research of methods for on-board data processing. Only recently, small datasets for multispectral sensors were made available in the works of [65]. However, works such as [66] show that real-time correction and preprocessing of the L0 hyperspectral data is feasible on low-cost and low-compute platforms. Similarly, the work of [67] uses partially processed data (so-called level 0.5) on-board of the EO-1 spacecraft. Furthermore, works such as [68] demonstrate that tiny machine learning models can serve as simulators for fast translation between processing levels (in that case L1C to L2A of Sentinel-2 data). For real world deployment, we’d suggest two approaches: either using even partial data processing pipelines on-board, or training deep models with simulated L0-like data.
For future research, we’d like to explore further steps to make the models more efficient. Our proposed changes to the Transformer based models model can be combined with pruning reported in [69, 40]. Specialised accelerators as those suggested by [70] could also be leveraged. Secondly, we’d also like to explore extension of our models with methods for uncertainty estimation. Methods such as model ensembles [71], or Monte Carlo Dropout [72] could be easily added to the proposed semantic segmentation models, as was previously demonstrated in [73]. Thirdly, we note that the existing archives of EMIT data (currently with more than 121k captures totaling more than 244 TB of data) offer unique opportunities to train large general foundation models on hyperspectral data, where one of the example tested tasks could also be methane detection. Models and datasets proposed in this paper could likely be leveraged for general learning tasks (in the direction outlined by our mineral identification task), but also for other downstream applications highlighted by the future planned missions such as NASA’s SBG [60] or ESA’s CHIME [74].
Finally, we are releasing the three newly created datasets at https://huggingface.co/previtus split into several repositories - OxHyperRealCH4, OxHyperSyntheticCH4 and OxHyperMinerals. The code and the pre-trained models are on Github at https://github.com/previtus/HyperspectralViTs.